In the short-run, capital is fixed at K. Find the short-run average total cost function when a firms production function is Q(L, K) 4. 1/2)2 L172-K 1 2 (r-w)K
Answer #1

Total cost = wL + rK(bar)

SRATC = (wL + rK(bar))/Q

Q0.5 = L 0.5 + K0.5

after rearranging and replacing K = K(bar):

L = (Q0.5 - K(bar)0.5)2

(replace 'L' in the first equation)

SRATC = (Q0.5 - K(bar)0.5)2 x w/Q + rK(bar)/Q

option b